# Chaos propagation in genetic algorithms: An optimal transport approach

### v1

@@ 3.1 Two particle systems
Note that $\mathbb{P}(\mathbf{j}(\mathbf{w},\alpha)=j)=w^{j}$ if $\alpha\sim\textup{Unif}[0,N)$. The particles evolution described by Algorithm 1 can then be written compactly as
@@ 3.2 Coupling of parents
For $p\geq 1$, let $\mathbb{W}_{p}$ be the Wasserstein distance with cost function $c(x,y)=|x-y|^{p}$, and let $f\in\mathcal{P}_{p}(\mathbb{R}^{d})$. Consider $\mathbf{x}=(x^{1},\dots,x^{N})\in(\mathbb{R}^{d})^{N}$ and weights $\mathbf{w}=(w^{1},\dots,w^{N})\in\textup{int}(\Delta_{N})$ such that $w^{i}>0$ and $\sum_{i=1}^{N}w^{i}=1$. Define the weighted empirical measure as $\mu_{\mathbf{x},\mathbf{w}}:=\sum_{i=1}^{N}w^{i}\,\delta_{x^{i}},$ and recall from (3.2) the index-map $\mathbf{j}(\mathbf{w},\cdot)$.
@@
with the following property: if $\alpha\sim\textup{Unif}[0,N)$, then the pair $\big(X_{f}^{*}(\mathbf{x},\mathbf{w},\alpha),\,x^{\mathbf{j}(\mathbf{w},\alpha)}\big)$ is an optimal coupling between $f$ and $\mu_{\mathbf{x},\mathbf{w}}$, namely
$\mathbb{E}\Big|X_{f}^{*}(\mathbf{x},\mathbf{w},\alpha)-x^{\mathbf{j}(\mathbf{w},\alpha)}\Big|^{p}=\mathbb{W}_{p}^{p}\big(f,\mu_{\mathbf{x},\mathbf{w}}\big).$
@@
Let $\pi_{\mathbf{x},\mathbf{w}}\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ be an optimal transport plan between $f$ and $\mu_{\mathbf{x},\mathbf{w}}$. Thanks to a measurable selection result, see, for instance, [33, Corollary 5.22], there exists a measurable mapping
@@
We note that $G^{i}$ is a probability kernel from $(\mathbb{R}^{d})^{N}\times\textup{int}(\Delta_{N})$ into $\mathbb{R}^{d}$, thanks to the measurability of $(\mathbf{x},\mathbf{w})\mapsto\pi_{\mathbf{x},\mathbf{w}}$, and so there exists $g^{i}=g^{i}(\mathbf{x},\mathbf{w},\beta)$ such that $\textup{Law}(g^{i}(\mathbf{x},\mathbf{w},\beta))=G^{i}(\mathbf{x},\mathbf{w},\cdot)$ if $\beta\sim\textup{Unif}[0,1)$, see [25, Lemma 4.22]. This procedure is called randomization of $G^{i}(\mathbf{x},\mathbf{w},\cdot)$.
Let $S_{0}:=0$ and $S_{i}:=\sum_{\ell=1}^{i}w^{\ell}$ for $i=1,\dots,N$. Define the rescaled variable
@@
Observe that, if $\alpha\sim\textup{Unif}[0,N)$, then $\mathbb{P}(\mathbf{j}(\mathbf{w},\alpha)=i)=w^{i}$ and, conditionally on $\{\mathbf{j}(\mathbf{w},\alpha)=i\}$, the random variable $\beta(\mathbf{w},\alpha)$ is uniform on $[0,1)$.
@@
To conclude, we need to show that $\big(X_{f}^{*}(\mathbf{x},\mathbf{w},\cdot),\,x^{\mathbf{j}(\mathbf{w},\cdot)}\big)$ has joint distribution $\pi_{\mathbf{x},\mathbf{w}}$ for $\alpha\sim\textup{Unif}[0,N)$. Take a Borel set $A\subseteq\mathbb{R}^{d}$ and $j\in\{1,\dots,N\}$, we have
@@
$\mathbb{E}\Big|X_{f}^{*}(\mathbf{x},\mathbf{w},\alpha)-x^{\mathbf{j}(\mathbf{w},\alpha)}\Big|^{p}=\int|x-y|^{p}\,\pi_{\mathbf{x},\mathbf{w}}(\mathrm{d}x,\mathrm{d}y)=\mathbb{W}_{p}^{p}\big(f,\mu_{\mathbf{x},\mathbf{w}}\big).$
@@ 3.3 Limit as $N\to\infty$
There exists a constant $\tilde{C}=\tilde{C}(q)>0$ such that for any $n\geq 0$ it holds $M^{1/q}_{q}(f_{n})\leq\tilde{C}e^{\tilde{C}\kappa_{\mathbf{F}}n\tau}(M^{1/q}_{q}(f_{0})+\sigma)$.
@@
thanks to Lemma 2.3 with $\phi(x)=|x|^{q}$. Hence $M_{q}(f_{n+1})\leq\left(1+\tau C\kappa_{\mathbf{F}}\right)M_{q}(f_{n})+\tau C\kappa_{\mathbf{F}}\sigma^{q}.$ Iterating the inequality yields
@@
Through the random variables $(\gamma^{i}_{n},\xi_{n}^{i},\tau_{n}^{i},\alpha^{i,1}_{n},\alpha^{i,2}_{n})$, we have coupled particle $i$ of the algorithm (3.3) with particle $i$ of the nonlinear system (3.4). By coupling optimality of $X^{*}_{n}$, we can track their distance over time as
$\mathbb{E}|X^{i}_{n+1}-\overline{X}_{n+1}^{i}|=(1-\tau)\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|$
$\qquad+\tau\mathbb{E}\left|(1-\gamma^{i}_{n})\odot\left(X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i,1}_{n})}_{n}-X^{*}_{n}(\alpha^{i,1}_{n})\right)+\gamma^{i}_{n}\odot\left(X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i,2}_{n})}_{n}-X^{*}_{n}(\alpha^{i,2}_{n})\right)\right|$
$\leq(1-\tau)\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|+\tau\left(\mathbb{E}\left|X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i,1}_{n})}_{n}-X^{*}_{n}(\alpha^{i,1}_{n})\right|+\mathbb{E}\left|X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i,2}_{n})}_{n}-X^{*}_{n}(\alpha^{i,2}_{n})\right|\right)$
$\leq(1-\tau)\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|+2\tau\mathbb{E}\mathbb{W}_{1}\Big(\tilde{f}_{n},\sum_{i=1}^{N}w_{n}^{i}\delta_{X^{i}_{n}}\Big)$ (3.5)
where we also used that $\mathbb{E}|\gamma^{i}_{n}\odot x|\leq|x|$, since $\textup{esssup}\,\gamma^{i}_{n}\leq 1$. Note that the additive noise $\xi_{n}^{i}$ disappeared as the same was used in both updates.
Recall the weights are such that $w^{i}_{n}\propto\mathbf{F}(X^{i}_{n})$. Since $f_{n}\in\mathcal{P}_{q}(\mathbb{R}^{d}),q\geq 1$ (Lemma 3.2), we also have $\tilde{f}_{n}\in\mathcal{P}_{q}(\mathbb{R}^{d})$ thanks to the assumption of $\mathbf{F}$. Therefore, any test function $\phi,\textup{Lip}(\phi)\leq 1$, also belongs to $L^{1}(\tilde{f}_{n})$. Let $\tilde{f}_{n}^{N}=\sum_{i}w^{i}_{n}\delta_{X^{i}_{n}}$, we apply Lemma 2.3 and obtain for any $\textup{Lip}(\phi)\leq 1$
$|\langle\phi,\tilde{f}_{n}-\tilde{f}^{N}_{n}\rangle|\leq\kappa_{\mathbf{F}}\langle|\phi|,{f}_{n}-{f}^{N}_{n}\rangle\leq\kappa_{\mathbf{F}}\sup_{\textup{Lip}(\psi)\leq 1}\langle\psi,{f}_{n}-{f}^{N}_{n}\rangle$
from which it follows by (2.1) $\mathbb{W}_{1}(\tilde{f}_{n},\tilde{f}_{n}^{N})\leq\kappa_{\mathbf{F}}\mathbb{W}_{1}(f_{n},f_{n}^{N})\,.$ By continuing the estimate (3.5), we then obtain
$\mathbb{E}|X^{i}_{n+1}-\overline{X}_{n+1}^{i}|$ $\leq(1-\tau)\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|+2\tau\kappa_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}(f_{n},f^{N}_{n})$
$\leq(1-\tau)\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|+2\tau\kappa_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}(f_{n},\overline{f}^{N}_{n})+2\tau\kappa_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}(\overline{f}^{N}_{n},f^{N}_{n})$
@@
Since $\mathbb{W}_{1}(f^{N}_{n},\overline{f}_{n}^{N})\leq(1/N)\sum_{i}|X^{i}_{n+1}-\overline{X}_{n+1}^{i}|$, when we sum the estimate among all $i=1,\dots,N$, we obtain
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}|X^{i}_{n+1}-\overline{X}_{n+1}^{i}|\leq(1+\tau(2\kappa_{\mathbf{F}}-1))\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|+2\tau\kappa_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}(f_{n},\overline{f}^{N}_{n})\,.$
It is crucial now to note that the term $\mathbb{E}\mathbb{W}_{1}(f_{n},\overline{f}^{N}_{n})$ is something that we can control via the concentration inequality provided by Lemma 2.2 since $\overline{f}_{n}^{N}$ is made of i.i.d $f_{n}$-distributed particles. In particular, in view of the moments estimate (Lemma 3.2), we have for constants $C(q,d,M_{q}(f_{0})^{1/q},\sigma)>0, C=C$$^{\prime}=C^{\prime}(q)>0$
$\mathbb{E}\mathbb{W}_{1}(f_{n},\overline{f}^{N}_{n})\leq Ce^{C\ \kappa_{\mathbf{F}}n\tau}\varepsilon_{1}(N)\,.$
We iterate now the estimates for $n\geq 0$. To simplify the notation, let
$E_{n}:=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}|X^{i}_{n}-\overline{X}_{n}^{i}|\,.$
Then the previous inequality implies
$E_{n+1}$ $\leq(1+2\tau\kappa_{\mathbf{F}})\,E_{n}+2\tau\kappa_{\mathbf{F}}\,\mathbb{E}\mathbb{W}_{1}(f_{n},\overline{f}_{n}^{N})$
$\leq(1+2\tau\kappa_{\mathbf{F}})\,E_{n}+C\,\tau\,\kappa_{\mathbf{F}}\,e^{C^{\prime}\kappa_{\mathbf{F}}n\tau}\varepsilon_{1}(N),$
for again, some constant $C=C(q,d,M_{q}(f_{0})^{1/q},\sigma)>0$. Iterating for $n\tau\leq T$, and using that $E_{0}=0$ yields the Grönwall-type inequality
$E_{n}$ $\leq(1+2\tau\kappa_{\mathbf{F}})^{n}E_{0}+C\,\tau\,\kappa_{\mathbf{F}}\,\varepsilon_{1}(N)\sum_{k=0}^{n-1}(1+2\tau\kappa_{\mathbf{F}})^{\,n-1-k}e^{C^{\prime}\kappa_{\mathbf{F}}k\tau}$
$\leq C\,\tau\,\kappa_{\mathbf{F}}\,\varepsilon_{1}(N)\sum_{k=0}^{n-1}e^{2\kappa_{\mathbf{F}}(n-1-k)\tau}\,e^{C^{\prime}\kappa_{\mathbf{F}}k\tau}$
$\leq C\,\tau\,\kappa_{\mathbf{F}}\,\varepsilon_{1}(N)\sum_{k=0}^{n-1}e^{2C^{\prime}\kappa_{\mathbf{F}}(n-1)\tau}\leq C\,\kappa_{\mathbf{F}}\,T\,e^{\tilde{C}^{\prime}\kappa_{\mathbf{F}}T}\,M_{q}(f_{0})^{1/q}\varepsilon_{1}(N).$
Finally, since $\mathbb{E}\mathbb{W}_{1}(f^{N}_{n},\overline{f}^{N}_{n})=\mathbb{E}\|f^{N}_{n}-\overline{f}_{n}^{N}\|_{\mathtt{KR}}\leq E_{n}$ by definition of the Wasserstein distance, we obtain the estimate
$\begin{split}\mathbb{E}\|f^{N}_{n}-f_{n}\|_{\mathtt{KR}}&\leq\mathbb{E}\|f^{N}_{n}-\overline{f}^{N}_{n}\|_{\mathtt{KR}}+\mathbb{E}\|\overline{f}^{N}_{n}-f_{n}\|_{\mathtt{KR}}\\ &\leq C(1+\kappa_{\mathbf{F}})Te^{\tilde{C}^{\prime}\kappa_{\mathbf{F}}T}\varepsilon_{1}(N)\end{split}$ (3.6)
[…]

### v2

@@ 3.1 Two particle systems
Note that $\mathbb{P}(\mathbf{j}(\mathbf{w},\alpha)=j)=w^{j}$ if $\alpha\sim\mathrm{Unif}[0,N)$. The particles evolution described by Algorithm 1 can then be written compactly as
@@ 3.2 Coupling of parents
For $p\geq 1$, let $\mathbb{W}^{c}_{p}$ be the Wasserstein distance with a lower semi-continuous cost function $c(x,y)$, and let $f\in\mathcal{P}(\mathbb{R}^{d})$ such that $\mathbb{W}_{p}^{c}(f,\delta_{0})<+\infty$. Consider $\mathbf{x}=(x^{1},\dots,x^{N})\in(\mathbb{R}^{d})^{N}$ and weights $\mathbf{w}=(w^{1},\dots,w^{N})\in\textup{int}(\Delta_{N})$ such that $w^{i}>0$ and $\sum_{i=1}^{N}w^{i}=1$. Define the weighted empirical measure as $\mu_{\mathbf{x},\mathbf{w}}:=\sum_{i=1}^{N}w^{i}\,\delta_{x^{i}},$ and recall from (3.2) the index-map $\mathbf{j}(\mathbf{w},\cdot)$.
@@
with the following property: if $\alpha\sim\mathrm{Unif}[0,N)$, then the pair $\big(X_{f}^{*}(\mathbf{x},\mathbf{w},\alpha),\,x^{\mathbf{j}(\mathbf{w},\alpha)}\big)$ is an optimal coupling between $f$ and $\mu_{\mathbf{x},\mathbf{w}}$, namely
$\mathbb{E}\left[c\Big(X_{f}^{*}(\mathbf{x},\mathbf{w},\alpha),x^{\mathbf{j}(\mathbf{w},\alpha)}\Big)^{p}\right]=\left(\mathbb{W}_{p}^{c}\big(f,\mu_{\mathbf{x},\mathbf{w}}\big)\right)^{p}.$
@@
Let $\pi_{\mathbf{x},\mathbf{w}}\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ be an optimal transport plan between $f$ and $\mu_{\mathbf{x},\mathbf{w}}$ (assumption $\mathbb{W}_{p}^{c}(f,\delta_{0})<+\infty$ ensures its existence). Thanks to a measurable selection result for lower semi-continuous costs, see, for instance, [39, Corollary 5.22], there exists a measurable mapping
@@
We note that $G^{i}$ is a probability kernel from $(\mathbb{R}^{d})^{N}\times\textup{int}(\Delta_{N})$ into $\mathbb{R}^{d}$, thanks to the measurability of $(\mathbf{x},\mathbf{w})\mapsto\pi_{\mathbf{x},\mathbf{w}}$, and so there exists $g^{i}=g^{i}(\mathbf{x},\mathbf{w},\beta)$ such that $\mathrm{Law}(g^{i}(\mathbf{x},\mathbf{w},\beta))=G^{i}(\mathbf{x},\mathbf{w},\cdot)$ if $\beta\sim\mathrm{Unif}[0,1)$, see [29, Lemma 4.22]. This procedure is called randomization of $G^{i}(\mathbf{x},\mathbf{w},\cdot)$.
Let $S_{0}:=0$ and $S_{i}:=\sum_{\ell=1}^{i}w^{\ell}$ for $i=1,\dots,N$. Since weights are strictly positive, define the rescaled variable
@@
Observe that, if $\alpha\sim\mathrm{Unif}[0,N)$, then $\mathbb{P}(\mathbf{j}(\mathbf{w},\alpha)=i)=w^{i}$ and, conditionally on $\{\mathbf{j}(\mathbf{w},\alpha)=i\}$, the random variable $\beta(\mathbf{w},\alpha)$ is uniform on $[0,1)$.
@@
To conclude, we need to show that $\big(X_{f}^{*}(\mathbf{x},\mathbf{w},\cdot),\,x^{\mathbf{j}(\mathbf{w},\cdot)}\big)$ has joint distribution $\pi_{\mathbf{x},\mathbf{w}}$ for $\alpha\sim\mathrm{Unif}[0,N)$. Take a Borel set $A\subseteq\mathbb{R}^{d}$ and $j\in\{1,\dots,N\}$, we have
@@
$\mathbb{E}\left[c\Big(X_{f}^{*}(\mathbf{x},\mathbf{w},\alpha),x^{\mathbf{j}(\mathbf{w},\alpha)}\Big)^{p}\right]=\int c(x,y)^{p}\,\pi_{\mathbf{x},\mathbf{w}}(\mathrm{d}x,\mathrm{d}y)=\left(\mathbb{W}_{p}^{c}\big(f,\mu_{\mathbf{x},\mathbf{w}}\big)\right)^{p}.$
@@ 3.3 Limit as $N\to\infty$
Let Assumption 1.2 hold for $\mathbf{F}$, and $(f_{n})_{N\in\mathbb{N}}$ be defined by (3.1) with $f_{0}\in\mathcal{P}_{q}(\mathbb{R}^{d})$. There exists a constant $\tilde{C}=\tilde{C}(q)>0$ such that for any $n\geq 0$ it holds $M^{1/q}_{q}(f_{n})\leq\tilde{C}e^{\tilde{C}\kappa_{\mathbf{F}}n\tau}(M^{1/q}_{q}(f_{0})+\sigma)$, with $\kappa_{\mathbf{F}}=\overline{\mathbf{F}}/\underline{\mathbf{F}}$.
@@
thanks to Assumption 1.2 on $\mathbf{F}$. Hence $M_{q}(f_{n+1})\leq\left(1+\tau C\kappa_{\mathbf{F}}\right)M_{q}(f_{n})+\tau C\kappa_{\mathbf{F}}\sigma^{q}.$ Iterating the inequality yields
@@
Through the random variables $(\gamma^{i}_{n},\xi_{n}^{i},\tau_{n}^{i},\alpha^{i,1}_{n},\alpha^{i,2}_{n})$, we have coupled particle $i$ of the algorithm (3.3) with particle $i$ of the nonlinear system (3.4). Recall the crossover-mutation update (1.1) is Lipschitz by with respect to $D^{\wedge}$} by Lemma 2.3. By coupling optimality of $X^{*}_{n}$, we can track their distance over time as
$\mathbb{E}{D^{\wedge}}(X^{i}_{n+1},\overline{X}_{n+1}^{i})$ $\leq(1-\tau)\mathbb{E}{D^{\wedge}}(X^{i}_{n},\overline{X}_{n}^{i})$
$\qquad+\tau\left(\mathbb{E}{D^{\wedge}}\left(X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i,1}_{n})}_{n},X^{*}_{n}(\alpha^{i,1}_{n})\right)+\mathbb{E}{D^{\wedge}}\left(X^{\mathbf{j}(\mathbf{w}_{n},\alpha^{i,2}_{n})}_{n},X^{*}_{n}(\alpha^{i,2}_{n})\right)\right)$
$\leq(1-\tau)\mathbb{E}{D^{\wedge}}(X^{i}_{n},\overline{X}_{n}^{i})+2\tau\mathbb{E}\mathbb{W}_{1}^{\wedge}\Big(\tilde{f}_{n},\sum_{i=1}^{N}w_{n}^{i}\delta_{X^{i}_{n}}\Big)$ (3.5)
Recall the weights are such that $w^{i}_{n}\propto\mathbf{F}(X^{i}_{n})$, and so $\tilde{f}_{n}^{N}:=\sum_{i}w^{i}_{n}\delta_{X^{i}_{n}}=\mathbf{F}f^{N}_{n}/\langle\mathbf{F},f^{N}_{n}\rangle$ is also a re-weighted probability measure like $\tilde{f}$. Thanks to Lemma 2.4 we then have $\mathbb{W}^{|\cdot|}_{1}(\tilde{f}_{n},\tilde{f}_{n}^{N})\leq C_{\mathbf{F}}\mathbb{W}^{|\cdot|}_{1}(f_{n},f_{n}^{N})\,.$ By continuing the estimate (3.5), we obtain
$\mathbb{E}{D^{\wedge}}(X^{i}_{n+1},\overline{X}_{n+1}^{i})$ $\leq(1-\tau)\mathbb{E}{D^{\wedge}}(X^{i}_{n},\overline{X}_{n}^{i})+2\tau C_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}^{\wedge}(f_{n},f^{N}_{n})$
$\leq(1-\tau)\mathbb{E}{D^{\wedge}}(X^{i}_{n},\overline{X}_{n}^{i})+2\tau C_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}^{\wedge}(f_{n},\overline{f}^{N}_{n})+2\tau C_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}^{\wedge}(\overline{f}^{N}_{n},f^{N}_{n})\,.$
@@
Since $\mathbb{W}_{1}^{\wedge}(f^{N}_{n},\overline{f}_{n}^{N})\leq(1/N)\sum_{i}{D^{\wedge}}(X^{i}_{n+1},\overline{X}_{n+1}^{i})$, when we sum the estimate among all $i=1,\dots,N$, we obtain
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}{D^{\wedge}}(X^{i}_{n+1},\overline{X}_{n+1}^{i})\leq(1+\tau(2C_{\mathbf{F}}-1))\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}{D^{\wedge}}(X^{i}_{n},\overline{X}_{n}^{i})+2\tau C_{\mathbf{F}}\mathbb{E}\mathbb{W}_{1}^{\wedge}(\overline{f}^{N}_{n},f_{n})\,.$
It is crucial now to note that the term $\mathbb{E}\mathbb{W}_{1}^{\wedge}(\overline{f}^{N}_{n},f_{n})$ is something that we can control thanks to $\mathbb{W}_{1}^{\wedge}\leq\mathbb{W}^{|\cdot|}_{1}$, and for $\mathbb{W}^{|\cdot|}_{1}$ we can use the concentration inequality provided by Lemma 2.2 since $\overline{f}_{n}^{N}$ is made of i.i.d $f_{n}$-distributed particles. In particular, in view of the moments estimate (Lemma 3.2), we have for constants $C(q,d,M_{q}(f_{0})^{1/q},\sigma)>0, C=C$$^{\prime}=C^{\prime}(q)>0$
$\mathbb{E}\mathbb{W}^{|\cdot|}_{1}(\overline{f}^{N}_{n},f_{n})\leq Ce^{C^{\prime}\kappa_{\mathbf{F}}n\tau}\varepsilon_{1}(N)\,.$
We iterate now the estimates for $n\geq 0$. To simplify the notation, let $E_{n}:=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}{D^{\wedge}}(X^{i}_{n},\overline{X}_{n}^{i})\,.$ Then the previous inequality implies
$E_{n+1}$ $\leq(1+2\tau C_{\mathbf{F}})\,E_{n}+2\tau C_{\mathbf{F}}\,\mathbb{E}\mathbb{W}_{1}^{\wedge}(f_{n},\overline{f}_{n}^{N})$
$\leq(1+2\tau C_{\mathbf{F}})\,E_{n}+C_{0}\,\tau\,C_{\mathbf{F}}\,e^{C^{\prime}C_{\mathbf{F}}n\tau}\varepsilon_{1}(N),$
for again, some constant $C_{0}=C_{0}(q,d,M_{q}(f_{0})^{1/q},\sigma)>0$. Iterating for $n\tau\leq T$, and using that $E_{0}=0$ yields the Grönwall-type inequality
$E_{n}$ $\leq(1+2\tau C_{\mathbf{F}})^{n}E_{0}+C_{0}\,\tau\,C_{\mathbf{F}}\,\varepsilon_{1}(N)\sum_{k=0}^{n-1}(1+2\tau C_{\mathbf{F}})^{\,n-1-k}e^{C^{\prime}C_{\mathbf{F}}k\tau}$
$\leq C_{0}\,\tau\,C_{\mathbf{F}}\,\varepsilon_{1}(N)\sum_{k=0}^{n-1}e^{2C_{\mathbf{F}}(n-1-k)\tau}\,e^{C^{\prime}C_{\mathbf{F}}k\tau}$
$\leq C_{0}\,\tau\,C_{\mathbf{F}}\,\varepsilon_{1}(N)\sum_{k=0}^{n-1}e^{C_{\mathbf{F}}\max(2,C^{\prime})(n-1)\tau}\leq C_{0}C_{\mathbf{F}}\,T\,e^{C_{\mathbf{F}}\max(2,C^{\prime})T}\,\varepsilon_{1}(N).$
Finally, since $\mathbb{E}\mathbb{W}_{1}^{\wedge}(f^{N}_{n},\overline{f}^{N}_{n})=\mathbb{E}\|f^{N}_{n}-\overline{f}_{n}^{N}\|_{\mathtt{BL}}\leq E_{n}$ by definition of the Wasserstein distance, we obtain the estimate
$\begin{split}\mathbb{E}\|f^{N}_{n}-f_{n}\|_{\mathtt{BL}}&\leq\mathbb{E}\|f^{N}_{n}-\overline{f}^{N}_{n}\|_{\mathtt{BL}}+\mathbb{E}\|\overline{f}^{N}_{n}-f_{n}\|_{\mathtt{BL}}\\ &\leq CTe^{C^{\prime}T}\varepsilon_{1}(N)\end{split}$ (3.6)
for some $C(\mathbf{F},q,d,M_{q}(f_{0})^{1/q},\sigma)>0, C=C$$^{\prime}=C^{\prime}(\mathbf{F},q)>0$.